Jack usually mows his yard in 7 hours. Jane can mow the same yard in 3 hours. How much time will it take to mow it together?
Answer by stanbon(75887)

Jack usually mows his yard in 7 hours. Jane can mow the same yard in 3 hours. How much time will it take to mow it together?
-------------------
Jack rate: 1/7 job/hr
Jane rate: 1/3 job/hr
----
Together rate: 1/x job/hr
-------------------------------
Equation:
1/7 + 1/3 = 1/x
Multiply thru by 21x to get:
3x + 7x = 21
10x = 21
x = 2.1 hrs (time to do the job together)
======================
Cheers,
Stan H.
==============